# A Diffused Background from Axion-like Particles in the Microwave Sky

### v1

@@ 8 Constraining the ALP coupling constant using ILC
The constraints on photon-ALP coupling constant $\mathrm{g_{a\gamma}}$ can be gauged using the constraints on the modified coupling constant A. We convert the constraints on A to the constraints on $\mathrm{g_{a\gamma}}$ using the fiducial values of the magnetic field and electron densities used to obtain A, i.e. $\mathrm{n_{0}=10^{-3}\,cm^{-3}}$ and $\mathrm{B_{0}=0.1\,\mu G}$. These are plotted in Fig.13 with the horizontal lines representing the 95% C.I. bounds for various detectors, with the allowed coupling strengths shaded. We expect an upper bound on ALP coupling strength of $\mathrm{g_{a\gamma}<1.929\times 10^{-12}\,GeV^{-1}}$ using SO and of about $\mathrm{g_{a\gamma}<1.119\times 10^{-12}\,GeV^{-1}}$ using CMB-S4 configurations for $\mathrm{z_{min}=1}$. ILC using CMB-HD can strongly constrain the coupling to $\mathrm{g_{a\gamma}<4.254\times 10^{-13}\,GeV^{-1}}$.
@@ 9 Conclusion
This background signal will depend on the cluster mass distribution at different redshifts, which are biased tracers of the dark matter halos. This is taken into account using the distribution of these halos and the correlation between different halos. The background will be an integrated effect of the signals from individual clusters of various masses at different redshifts. This diffused spectrum can be modelled using the distribution of halos of various masses at different redshifts. The two-halo or clustering component is low at high multipoles, but may dominate over the one-halo term or the Poissonian component for low ones (20 to 200), where it is itself dominated by the CMB and is difficult to probe. The one-halo or Poissonian component of the ALP power spectrum dominates at high multipoles and can be probed using the upcoming high resolution experiments (CMB-S4, CMB-HD) with low noise. The ALP background spectrum will depend on the astrophysical aspects like the electron density and magnetic field profiles in clusters (see Figs. 7). Also, it will increase with ALP coupling constant ($\mathrm{\propto g_{a\gamma}^{4}}$) (see Fig. 6) and the frequency of observation ($\mathrm{\propto\nu^{2}I_{cmb}(\nu)^{2}}$)(see Fig. 8).
@@
The background spectrum will increase as we lower $\mathrm{z_{min}}$, as low redshift clusters contribute significantly to the background spectrum (Fig. 4). With the upcoming experiments, clusters up to redshift z = 1 will be well resolvable, thus $\mathrm{z_{min}=1}$ is a conservative choice for the estimation of the ALP background spectrum. The power spectrum is almost independent of the choice of maximum redshift $\mathrm{z_{max}}$ after a certain redshift ($\sim 3.5$) as the clusters at very high redshifts contribute negligibly to the ALP background spectrum (Appendix B). The cluster mass range that will contribute to the background will also determine the strength of the diffused power spectrum (Fig.5). The ALP masses that may exist will also affect the background spectrum, with a weaker power spectrum for high mass ALPs.
@@
For a coupling constant of $\mathrm{10^{-12}\,GeV}$ and $\mathrm{z_{min}=1}$, with randomly generated cluster profiles for cluster masses $\mathrm{10^{13}-7\times 10^{15}\,M_{\odot}}$, the SNR is 4.36 in the 145 GHz band of CMB-S4, while it is around 93.87 in the 150 GHz band of CMB-HD (Fig. 9). Also, such a diffused signal would lead to RMS fluctuations of the order of $\mathrm{7.5\times 10^{-2}\,\mu K}$ for an ALP coupling constant of $\mathrm{g_{a\gamma}=10^{-12}\,GeV^{-1}}$ at 145 GHz. The frequency channels 90 - 160 GHz provide the best SNR for the ALP signal, owing to a decrease in foregrounds contamination and improved beam resolution.
@@
Techniques such as ILC (see Sec.7) can be used to mitigate the effect of foregrounds and CMB by using the spectral variation of the ALP signal (Fig.8). Using ILC for different detectors, we obtain the following bounds on the modified coupling constant (see Sec.8)- SO: A < 3.724; CMB-S4: A < 1.254; CMB-HD: A < 0.181 for the case of $\mathrm{z_{min}=1}$. The converted constraints on the coupling constant $\mathrm{g_{a\gamma}}$ can be gauged using the fiducial electron density and magnetic field strength values. The conversion provides the bounds of $\mathrm{g_{a\gamma}<1.783\times 10^{-12}\,GeV^{-1}}$ with SO and $\mathrm{g_{a\gamma}<1.056\times 10^{-12}\,GeV^{-1}}$ with CMB-S4 configuration for $\mathrm{z_{min}=0.5}$. CMB-HD can provide much tighter bounds of $\mathrm{g_{a\gamma}<3.912\times 10^{-13}\,GeV^{-1}}$. For $\mathrm{z_{min}=1}$, the bounds go as: SO - $\mathrm{g_{a\gamma}<1.929\times 10^{-12}\,GeV^{-1}}$ ; CMB-S4 - $\mathrm{g_{a\gamma}<1.119\times 10^{-12}\,GeV^{-1}}$; and CMB-HD - $\mathrm{g_{a\gamma}<4.254\times 10^{-13}\,GeV^{-1}}$.
@@
The template matching of foregrounds can tighten the constraints on coupling constant as shown in Appendix D. The shape of the foregrounds power spectrum and their violation of statistical isotropy can be used to reduce further contamination of foregrounds. This can help in achieving a higher SNR and better constraints. The galaxy clusters, with their intracluster medium (ICM) and galaxies make sites for numerous processes such as the SZ effects, lensing, CIB, synchrotron, etc. The ALP power spectrum can also be cross-correlated with diffused spectrums from these phenomena and the large scale structure to obtain better constraints on the ALP coupling constant and masses.
@@
. The photon-ALP conversion is confined to galaxy clusters which occupy small angular scales on the sky. For a particular ALP mass range, the signal is formed in a spherical shell around the cluster center for a spherically symmetric galaxy cluster. These shells are projected as a disk in 2d. This signal disk increases the ALP power spectrum at low angular scales. It is the higher multipoles that contain ALP signatures in the CMB spectrum as multipoles vary inversely with the angular scales $\mathrm{\Delta\theta\propto\frac{180^{o}}{\ell}}$.
@@
The astrophysics of galaxy clusters affect their electron density and magnetic field profiles, and hence the ALP distortion spectrum $|\alpha_{\ell}|^{2}$. For clusters with stochastic and turbulent electron density and magnetic field profiles, the polarization information of the ALP distortion will be lost as depolarization will be caused due to multiple conversions. For such cases, hydrodynamical simulations or observations can be used that fit the data well. We consider the case where this turbulence is negligible and model the electron density and magnetic profiles using smooth profiles, which fit the profiles for resolved clusters well.
@@
$\mathrm{n_{e}^{2}=Z\left[n_{0}^{2}\frac{(r/r_{c1})^{-\alpha}}{(1+r^{2}/r_{c1}^{2})^{3\beta-\alpha/2}}\frac{1}{(1+r^{\gamma}/r_{s}^{\gamma})^{\epsilon/\gamma}}+\frac{n_{02}^{2}}{(1+r^{2}/r_{c2}^{2})^{3\beta_{2}}}\right].}$ (A.1)
@@
$\mathrm{B(r)=B_{0}r^{-s}}.$ (A.2)
The contribution to the Q and U polarization stokes modes depends on the magnetic field direction at the conversion location. The magnetic field direction has been assumed to be uniformly randomly oriented. The magnetic field coherence scale is assumed to be greater than the beam of the instrument, otherwise the signal will be depolarized by the turbulent fields. The power spectrum $|\alpha_{\ell}|^{2}$ is that of the combined power in the two maps for which the polarized intensity $\mathrm{I_{pol}}$ is given as:
$\mathrm{I_{pol}=\sqrt{I_{Q}^{2}+I_{U}^{2}}}.$ (A.3)
@@
The power spectrum $|\alpha_{\ell}|^{2}$ will also vary for the mass range of ALPs being considered. ALPs of masses in a particular sub-range of the mass range $10^{-15}-10^{-11}$ eV may be assumed to be forming in the galaxy clusters if the resonant condition is satisfied ($\mathrm{m_{a}=m_{\gamma}}$). For our analysis, we assume that if ALPs exist, all ALPs of masses in the range $10^{-15}-10^{-11}$ eV will be produced during conversion. Also, we assign a uniform coupling constant of $\mathrm{10^{-12}\,GeV^{-1}}$ for all ALP masses in this range.
@@
The features of the ALP power spectrum for a single cluster $|\alpha_{\ell}|^{2}$ depend on the host cluster and its properties that characterize the ALP distortion signal. The variation of the ALP power spectrum from a single cluster $|\alpha_{\ell}|^{2}$ with the photon-ALP coupling constant and the frequency of observation is shown in Fig. 14. The ALP distortion signal scales as $\mathrm{g^{2}_{a,\gamma}}$ and $\mathrm{\nu I_{cmb}(\nu)}$ (Eq. 3.4) for frequencies of observation in the microwave and radio bands. Thus, the power spectrum $|\alpha_{\ell}|^{2}$ varies as $\mathrm{g^{4}_{a,\gamma}}$ and $\mathrm{\nu^{2}I_{cmb}(\nu)^{2}}$, where $\mathrm{I_{cmb}(\nu)}$ is the Planck black-body function (see Eq.3.7). Also shown in Fig.14 is the CMB power spectrum at small scales around the cluster region.
The probability of conversion at a location is inversely proportional to the gradient of electron density. An increase in the electron density ($\mathrm{n_{e}\propto n_{0}}$) (Fig. 15(a)) reduces the ALP power spectrum strength. The electron density steepness parameter($\beta_{1}$) and scaling radius ($\mathrm{r_{s}}$) control the rate at which the electron density varies with distance from cluster center. The spectrum increases at high multipoles with decreasing scaling radius and increasing steepness parameter as the electron density reduces in the outer regions of the cluster (Fig. 15(b) and 15(c)). Also, the spectrum is directly proportional to the square of the magnetic field. So, a steep decrease in magnetic field (parameterized by "s") leads to a weaker spectrum as the magnetic field reduces in the outer regions of the cluster where the conversion probability is high (Fig. 15(d)).
The effect of change in profiles on $|\alpha_{\ell}|^{2}$ is dominated by the the effect in the production of low mass ALPs. This is because the strength of the ALP distortion signal is generally high for low mass ALPs, owing to higher conversion probabilities in the outer regions of the cluster with low electron densities ($\mathrm{\propto\frac{1}{|\nabla n_{e}(r)|}}$). This, in turn shows up in the effect on $|\alpha|^{2}$. If ALPs of masses in only a sub-range of what we have considered ($10^{-15}-10^{-11}$ eV) are formed, the dependence of the ALP background spectrum on these profile parameters can change, especially in the case of formation of only high mass ALPs. This is considered in a separate work [15].
@@
The shape of $|\alpha_{\ell}|^{2}$’s at high multipoles depends on the magnetic field orientation that characterizes the polarization of the ALP signal. If the magnetic field orientation is uniformly random, as in our analysis, $|\alpha_{\ell}|^{2}$ scales as $\ell^{0}$ ($\mathrm{D_{\ell}}$ as $\ell^{2}$) at high multipoles.
@@ Appendix B ALP background spectrum Variation with maximum redshift
In principle, clusters from very high redshifts will contribute to the diffused spectrum. But this contribution will be very low and the power spectrum is nearly independent of the maximum redshift $\mathrm{z_{max}}$ up to which we consider the distribution of clusters as can be seen in Fig. 16 as the spectrums at various redshifts overlap. The halo mass function decreases significantly with increasing redshift and hence, the number of clusters contributing at very high redshifts ($\mathrm{z_{max}}>3.5$) to the one and two-halo terms decreases significantly. Also, as a polarized photon travels from high redshifts, it may get depolarized due to multiple scatterings in galaxy clusters. Hence, it is mainly the polarized distortion signals from low redshift clusters that remain intact and contribute to the background ALP spectrum. This explains the choice of $\mathrm{z_{max}}$ of 7 for our analysis as it reduces computation time.
@@ Appendix C Power spectrum estimation from a CMB map
$\mathrm{\Delta^{net}=\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}a_{\ell m}Y_{\ell m}(\theta,\phi)}.$ (C.1)
@@
$\mathrm{\langle a_{\ell m}^{net*}a_{\ell^{\prime}m^{\prime}}^{net}\rangle=C_{\ell}^{net}\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}}.$ (C.2)
@@
$\mathrm{\Delta^{net}=\Delta^{cmb}+\Delta^{ax}+\Delta^{fg}}.$ (C.3)
@@
$\mathrm{C_{\ell}^{net}=\langle a_{\ell m}^{ax*}a_{\ell m}^{ax}\rangle=\sum_{j}\sum_{i}\langle a_{\ell m}^{i*}a_{\ell m}^{j}\rangle=\langle a_{\ell m}^{cmb*}a_{\ell m}^{cmb}\rangle+\langle a_{\ell m}^{ax*}a_{\ell m}^{ax}\rangle+\langle a_{\ell m}^{fg*}a_{\ell m}^{fg}\rangle}.$ (C.4)
@@
$\mathrm{a_{\ell m}^{obs}=B_{\ell}(a_{\ell m}^{cmb}+a_{\ell m}^{ax}+a_{\ell m}^{fg})+\eta_{\ell m}},$ (C.5)
where $\mathrm{B_{\ell}=exp(-\ell(\ell+1)\theta_{beam}^{2}/2)}$ and $\eta_{\ell m}$ are the fourier coefficients introduced due to instrumental noise.
The coefficients $\mathrm{a_{\ell m}^{i}}$s are assumed to follow a Gaussian distribution with mean zero and variance given by the corresponding $\mathrm{C_{\ell}^{i}}$s, i.e.,
$\mathrm{P(a_{\ell m}^{i}|C_{\ell}^{i})=\frac{1}{\sqrt{2\pi C_{\ell}^{i}}}exp\left(-\frac{|a_{\ell m}^{i}|^{2}}{2C_{\ell}^{i}}\right)}.$ (C.6)
@@
$\mathrm{\langle\eta_{\ell m}*\eta_{\ell^{\prime}m^{\prime}}\rangle=N_{\ell}^{i}\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}}.$ (C.7)
The $\mathrm{a_{\ell m}^{obs}}$s are assumed to follow a Gaussian distribution with mean $\mathrm{B_{\ell}\sum_{i}a_{\ell m}^{i}}$ and variance $\mathrm{N_{\ell}}$ given as:
$\mathrm{P(a_{{\ell}m}^{obs}|\{a_{\ell m}^{i}\})=\frac{1}{\sqrt{2\pi N_{\ell}}}exp\left(-\frac{|a_{\ell m}^{obs}-\sum_{i}B_{\ell}a_{\ell m}^{i}|^{2}}{2N_{\ell}}\right)}.$ (C.8)
@@
$\begin{split}\mathrm{P(a_{\ell m}^{obs}|\{C_{\ell}^{i}\})=\prod_{m=-\ell}^{\ell}\int\int\int\prod_{i}d\,a_{\ell m}^{i}P(a_{\ell m}^{obs}|a_{\ell m}^{i})P(a_{\ell m}^{i}|C_{\ell}^{i})},\\ \mathrm{=[2\pi(B_{\ell}^{2}\sum_{i}C_{\ell}^{i}+N_{\ell})]^{-(2\ell+1)/2}exp\left[\sum_{m=-\ell}^{\ell}-\frac{|a_{\ell m}^{obs}|^{2}}{2(\sum_{i}C_{\ell}^{i}B_{\ell}^{2}+N_{\ell})}\right]}.\end{split}$ (C.9)
We find the maximum likelihood estimator for the power spectrum of the component i by differentiating with respect to $\mathrm{C_{\ell}^{ax}}$ and setting equal to zero. It can be written as:
$\mathrm{\tilde{C_{\ell}^{i}}=B_{\ell}^{-2}\left[\frac{1}{2\ell+1}\sum_{m=-\ell}^{\ell}|a_{\ell m}^{obs}|^{2}-N_{\ell}\right]-\sum_{j\neq i}C_{\ell}^{j}}.$ (C.10)

### v3

@@ 8 Constraining the ALP coupling constant using ILC
The constraints on photon-ALP coupling constant $g_{a\gamma}$ can be gauged using the constraints on the modified coupling constant $A$. We convert the constraints on $A$ to the constraints on $g_{a\gamma}$ using the fiducial values of the magnetic field and electron densities used to obtain $A$, i.e. $n_{0}=10^{-3}\,\mathrm{cm^{-3}}$ and $B_{0}=0.1\,\mathrm{\mu G}$. These are plotted in Fig.13 with the horizontal lines representing the 95% C.I. bounds for various detectors, with the allowed coupling strengths shaded. We expect an upper bound on ALP coupling strength of $g_{a\gamma}<1.929\times 10^{-12}\,\mathrm{GeV^{-1}}$ using SO and of about $g_{a\gamma}<1.119\times 10^{-12}\,\mathrm{GeV^{-1}}$ using CMB-S4 configurations for $z_{\mathrm{min}}=1$. ILC using CMB-HD can strongly constrain the coupling to $g_{a\gamma}<4.254\times 10^{-13}\,\mathrm{GeV^{-1}}$.
@@ 9 Conclusion
This background signal will depend on the cluster mass distribution at different redshifts, which are biased tracers of the dark matter halos. This is taken into account using the distribution of these halos and the correlation between different halos. The background will be an integrated effect of the signals from individual clusters of various masses at different redshifts. This diffused spectrum can be modelled using the distribution of halos of various masses at different redshifts. The two-halo or clustering component is low at high multipoles, but may dominate over the one-halo term or the Poissonian component for low ones (20 to 200), where it is itself dominated by the CMB and is difficult to probe. The one-halo or Poissonian component of the ALP power spectrum dominates at high multipoles and can be probed using the upcoming high resolution experiments (CMB-S4, CMB-HD) with low noise. The ALP background spectrum will depend on the astrophysical aspects like the electron density and magnetic field profiles in clusters (see Figs. 7). Also, it will increase with ALP coupling constant ($\propto g_{a\gamma}^{4}$) (see Fig. 6) and the frequency of observation ($\propto\nu^{2}I_{\mathrm{cmb}}(\nu)^{2}$)(see Fig. 8).
@@
The background spectrum will increase as we lower $z_{\mathrm{min}}$, as low redshift clusters contribute significantly to the background spectrum (Fig. 4). With the upcoming experiments, clusters up to redshift $z=1$ will be well resolvable, thus $z_{\mathrm{min}}=1$ is a conservative choice for the estimation of the ALP background spectrum. The power spectrum is almost independent of the choice of maximum redshift $z_{\mathrm{max}}$ after a certain redshift ($\sim 3.5$) as the clusters at very high redshifts contribute negligibly to the ALP background spectrum (Appendix B). The cluster mass range that will contribute to the background will also determine the strength of the diffused power spectrum (Fig.5). The ALP masses that may exist will also affect the background spectrum, with a weaker power spectrum for high mass ALPs.
@@
For a coupling constant of $10^{-12}\,\mathrm{GeV^{-1}}$ and $z_{\mathrm{min}}=1$, with randomly generated cluster profiles for cluster masses $\mathrm{10^{13}-7\times 10^{15}\,M_{\odot}}$, the SNR is 4.36 in the 145 GHz band of CMB-S4, while it is around 93.87 in the 150 GHz band of CMB-HD (Fig. 9). Also, such a diffused signal would lead to RMS fluctuations of the order of $\mathrm{7.5\times 10^{-2}\,\mu K}$ for an ALP coupling constant of $g_{a\gamma}=10^{-12}\,\mathrm{GeV^{-1}}$ at 145 GHz. The frequency channels 90 - 160 GHz provide the best SNR for the ALP signal, owing to a decrease in foregrounds contamination and improved beam resolution.
@@
Techniques such as ILC (see Sec.7) can be used to mitigate the effect of foregrounds and CMB by using the spectral variation of the ALP signal (Fig.8). Using ILC for different detectors, we obtain the following bounds on the modified coupling constant $A$ (see Sec.8):
$\mathrm{SO}:A<3.724;$
$\mathrm{CMB-S4}:A<1.254;$
$\mathrm{CMB-HD}:A<0.181,$
for the case of $z_{\mathrm{min}}=1$. The converted constraints on the coupling constant $g_{a\gamma}$ can be gauged using the fiducial electron density and magnetic field strength values. The conversion provides the bounds of $g_{a\gamma}<1.783\times 10^{-12}\,\mathrm{GeV^{-1}}$ with SO and $g_{a\gamma}<1.056\times 10^{-12}\,\mathrm{GeV^{-1}}$ with CMB-S4 configuration for $z_{\mathrm{min}}=0.5$. CMB-HD can provide much tighter bounds of $g_{a\gamma}<3.912\times 10^{-13}\,\mathrm{GeV^{-1}}$. For $z_{\mathrm{min}}=1$, the bounds go as:
$\mathrm{SO}:g_{a\gamma}<1.929\times 10^{-12}\,\mathrm{GeV^{-1}};$
$\mathrm{CMB-S4}:g_{a\gamma}<1.119\times 10^{-12}\,\mathrm{GeV^{-1}};$
$\mathrm{CMB-HD}:g_{a\gamma}<4.254\times 10^{-13}\,\mathrm{GeV^{-1}}.$
@@
The template matching of foregrounds can tighten the constraints on coupling constant as shown in Appendix D. The shape of the foregrounds power spectrum and their violation of statistical isotropy can be used to reduce further contamination of foregrounds. This can help in achieving a higher SNR and better constraints. The galaxy clusters, with their intracluster medium (ICM) and galaxies make sites for numerous processes such as the SZ effects, lensing, CIB, synchrotron, etc. The ALP power spectrum can also be cross-correlated with diffused spectra from these phenomena and the large-scale structure to obtain better constraints on the ALP coupling constant and masses [61].
@@
. The photon-ALP conversion is confined to galaxy clusters which occupy small angular scales on the sky. For a particular ALP mass range, the signal is formed in a spherical shell around the cluster center for a spherically symmetric galaxy cluster. These shells are projected as a disk in 2d. This signal disk increases the ALP power spectrum at low angular scales. It is the higher multipoles that contain ALP signatures in the CMB spectrum as multipoles vary inversely with the angular scales $\mathrm{\Delta\theta\propto 180^{o}/\ell}$.
@@
The astrophysics of galaxy clusters affects their electron density and magnetic field profiles, and hence the ALP distortion spectrum $|\alpha_{\ell}|^{2}$. For clusters with stochastic and turbulent electron density and magnetic field profiles, the polarization information of the ALP distortion will be lost as depolarization will be caused due to multiple conversions. For such cases, hydrodynamical simulations or observations can be used that fit the data well. We consider the case where this turbulence is negligible and model the electron density and magnetic profiles using smooth profiles, which fit the profiles for resolved clusters well.
@@
$n_{e}^{2}=Z\left[n_{0}^{2}\frac{(r/r_{c1})^{-\alpha}}{(1+r^{2}/r_{c1}^{2})^{3\beta-\alpha/2}}\frac{1}{(1+r^{\gamma}/r_{s}^{\gamma})^{\epsilon/\gamma}}+\frac{n_{02}^{2}}{(1+r^{2}/r_{c2}^{2})^{3\beta_{2}}}\right].$ (A.1)
@@
$B(r)=B_{0}r^{-s}.$ (A.2)
The contribution to the Q and U polarization stokes modes depends on the magnetic field direction at the conversion location. The magnetic field direction has been assumed to be uniformly randomly oriented. The magnetic field coherence scale is assumed to be greater than the beam of the instrument, otherwise the signal will be depolarized by the turbulent fields. The power spectrum $|\alpha_{\ell}|^{2}$ is that of the combined power in the two maps for which the polarized intensity $I_{\mathrm{pol}}$ is given as:
$I_{\mathrm{pol}}=\sqrt{I_{\mathrm{Q}}^{2}+I_{\mathrm{U}}^{2}}.$ (A.3)
@@
The power spectrum $|\alpha_{\ell}|^{2}$ will also vary for the mass range of ALPs being considered. ALPs of masses in a particular sub-range of the mass range $10^{-15}-10^{-11}$ eV may be assumed to be forming in the galaxy clusters if the resonant condition is satisfied ($m_{a}=m_{\gamma}$). For our analysis, we assume that if ALPs exist, all ALPs of masses in the range $10^{-15}-10^{-11}$ eV will be produced during conversion. Also, we assign a uniform coupling constant of $\mathrm{10^{-12}\,GeV^{-1}}$ for all ALP masses in this range.
@@
The features of the ALP power spectrum for a single cluster $|\alpha_{\ell}|^{2}$ depend on the host cluster and its properties that characterize the ALP distortion signal. The variation of the ALP power spectrum from a single cluster $|\alpha_{\ell}|^{2}$ with the photon-ALP coupling constant and the frequency of observation is shown in Fig. 14. The ALP distortion signal scales as $g^{2}_{a\gamma}$ and $\nu I_{\mathrm{cmb}}(\nu)$ (Eq. 3.5) for frequencies of observation in the microwave and radio bands. Thus, the power spectrum $|\alpha_{\ell}|^{2}$ varies as $g^{4}_{a\gamma}$ and $\nu^{2}I_{\mathrm{cmb}}(\nu)^{2}$, where $I_{\mathrm{cmb}}(\nu)$ is the Planck black-body function (see Eq.3.8). Also shown in Fig.14 is the CMB power spectrum at small scales around the cluster region.
The probability of conversion at a location is inversely proportional to the gradient of electron density. An increase in the electron density ($n_{e}\propto n_{0}$) (Fig. 15(a)) reduces the ALP power spectrum strength. The electron density steepness parameter($\beta_{1}$) and scaling radius ($r_{s}$) control the rate at which the electron density varies with distance from the cluster center. The spectrum increases at high multipoles with decreasing scaling radius and increasing steepness parameter as the electron density reduces in the outer regions of the cluster (Fig. 15(b) and 15(c)). Also, the spectrum is directly proportional to the square of the magnetic field. So, a steep decrease in magnetic field (parameterized by "$s$") leads to a weaker spectrum as the magnetic field reduces in the outer regions of the cluster where the conversion probability is high (Fig. 15(d)).
The effect of change in profiles on $|\alpha_{\ell}|^{2}$ is dominated by the effect in the production of low mass ALPs. This is because the strength of the ALP distortion signal is generally high for low mass ALPs, owing to higher conversion probabilities in the outer regions of the cluster with low electron densities ($\propto|\nabla n_{e}(r)|^{-1}$). This, in turn, shows up in the effect on $|\alpha_{\ell}|^{2}$. If ALPs of masses in only a sub-range of what we have considered ($10^{-15}-10^{-11}$ eV) are formed, the dependence of the ALP background spectrum on these profile parameters can change, especially in the case of formation of only high mass ALPs. This is considered in a separate work [34].
@@
The shape of $|\alpha_{\ell}|^{2}$’s at high multipoles depends on the magnetic field orientation that characterizes the polarization of the ALP signal. If the magnetic field orientation is uniformly random, as in our analysis, $|\alpha_{\ell}|^{2}$ scales as $\ell^{0}$ ($D_{\ell}$ as $\ell^{2}$) at high multipoles.
@@ Appendix B ALP background spectrum Variation with maximum redshift
In principle, clusters from very high redshifts will contribute to the diffused spectrum. But this contribution will be very low and the power spectrum is nearly independent of the maximum redshift $z_{\mathrm{max}}$ up to which we consider the distribution of clusters as can be seen in Fig. 16 as the spectra at various redshifts overlap. The halo mass function decreases significantly with increasing redshift and hence, the number of clusters contributing at very high redshifts ($z_{\mathrm{max}}>3.5$) to the one and two-halo terms decreases significantly. Also, as a polarized photon travels from high redshifts, it may get depolarized due to multiple scatterings in galaxy clusters. Hence, it is mainly the polarized distortion signals from low redshift clusters that remain intact and contribute to the background ALP spectrum. This explains the choice of $z_{\mathrm{max}}$ of 7 for our analysis as it reduces computation time.
@@ Appendix C Power spectrum estimation from a CMB map
$\Delta^{\mathrm{net}}=\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}a_{\ell m}Y_{\ell m}(\theta,\phi).$ (C.1)
@@
$\langle a_{\ell m}^{\mathrm{net*}}a_{\ell^{\prime}m^{\prime}}^{\mathrm{net}}\rangle=C_{\ell}^{\mathrm{net}}\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}.$ (C.2)
@@
$\Delta^{\mathrm{net}}=\Delta^{\mathrm{cmb}}+\Delta^{\mathrm{ax}}+\Delta^{\mathrm{fg}}.$ (C.3)
@@
$C_{\ell}^{\mathrm{net}}=\langle a_{\ell m}^{\mathrm{ax*}}a_{\ell m}^{\mathrm{ax}}\rangle=\sum_{j}\sum_{i}\langle a_{\ell m}^{i*}a_{\ell m}^{j}\rangle=\langle a_{\ell m}^{\mathrm{cmb*}}a_{\ell m}^{\mathrm{cmb}}\rangle+\langle a_{\ell m}^{\mathrm{ax*}}a_{\ell m}^{\mathrm{ax}}\rangle+\langle a_{\ell m}^{\mathrm{fg*}}a_{\ell m}^{\mathrm{fg}}\rangle.$ (C.4)
@@
$a_{\ell m}^{\mathrm{obs}}=B_{\ell}(a_{\ell m}^{\mathrm{cmb}}+a_{\ell m}^{\mathrm{ax}}+a_{\ell m}^{\mathrm{fg}})+\eta_{\ell m},$ (C.5)
where $B_{\ell}=\exp(-\ell(\ell+1)\theta_{\mathrm{beam}}^{2}/2)$ and $\eta_{\ell m}$ are the fourier coefficients introduced due to instrumental noise.
The coefficients $a_{\ell m}^{i}$’s are assumed to follow a Gaussian distribution with mean zero and variance given by the corresponding $C_{\ell}^{i}$’s, i.e.,
$P(a_{\ell m}^{i}|C_{\ell}^{i})=\frac{1}{\sqrt{2\pi C_{\ell}^{i}}}\exp\left(-\frac{|a_{\ell m}^{i}|^{2}}{2C_{\ell}^{i}}\right).$ (C.6)
@@
$\langle\eta_{\ell m}*\eta_{\ell^{\prime}m^{\prime}}\rangle=N_{\ell}^{i}\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}.$ (C.7)
The $a_{\ell m}^{\mathrm{obs}}$s are assumed to follow a Gaussian distribution with mean $B_{\ell}\sum_{i}a_{\ell m}^{i}$ and variance $N_{\ell}$ given as:
$P(a_{{\ell}m}^{\mathrm{obs}}|\{a_{\ell m}^{i}\})=\frac{1}{\sqrt{2\pi N_{\ell}}}\exp\left(-\frac{|a_{\ell m}^{\mathrm{obs}}-\sum_{i}B_{\ell}a_{\ell m}^{i}|^{2}}{2N_{\ell}}\right).$ (C.8)
@@
$\begin{split}P(a_{\ell m}^{\mathrm{obs}}|\{C_{\ell}^{i}\})=\prod_{m=-\ell}^{\ell}\int\int\int\prod_{i}\mathrm{d}\,a_{\ell m}^{i}P(a_{\ell m}^{\mathrm{obs}}|a_{\ell m}^{i})P(a_{\ell m}^{i}|C_{\ell}^{i}),\\ =[2\pi(B_{\ell}^{2}\sum_{i}C_{\ell}^{i}+N_{\ell})]^{-(2\ell+1)/2}\exp\left[\sum_{m=-\ell}^{\ell}-\frac{|a_{\ell m}^{\mathrm{obs}}|^{2}}{2(\sum_{i}C_{\ell}^{i}B_{\ell}^{2}+N_{\ell})}\right].\end{split}$ (C.9)
We find the maximum likelihood estimator for the power spectrum of the component $i$ by differentiating with respect to $C_{\ell}^{i}$ and setting equal to zero. It can be written as:
$\tilde{C_{\ell}^{i}}=B_{\ell}^{-2}\left[\frac{1}{2\ell+1}\sum_{m=-\ell}^{\ell}|a_{\ell m}^{\mathrm{obs}}|^{2}-N_{\ell}\right]-\sum_{j\neq i}C_{\ell}^{j}.$ (C.10)